OpenStudy (anonymous):

Can someone please help me Find value of cos^-1(-sqrt2/2)

OpenStudy (anonymous):

\[\cos^{-1} \left( -\sqrt{2}/2 \right)\]

OpenStudy (anonymous):

135

OpenStudy (ybarrap):